kathi and robert had a pottery stand at the annual skipack craft fair they sold some of theirpottery at the original price of $9.50 for each piece but later decreased the price of each piece by $2 if they sold all 90 pieces and took in $705 find how many pieces they sold at the original price and how many pieces they sold at the reduced price?
Answer by ankor@dixie-net.com(22740)

Kathi and Robert had a pottery stand at the annual skip jack craft fair they sold some of their pottery at the original price of $9.50 for each piece but later decreased the price of each piece by $2
if they sold all 90 pieces and took in $705 find how many pieces they sold at the original price and how many pieces they sold at the reduced price?
:
let n = number sold at original price
then
(90-n) = number sold at reduced price
:
9.50n + 7.50(90-n) = 705
9.50n + 675 - 7.50n = 705
9.5n - 7.5n = 705 - 675
2n = 30
n = 15 units at the original price
then
90-15 = 75 units at the reduced price
:
:
Check this find the total$
15(9.50) + 75(7.50) =
142.50 + 562.50 = 705